OpenStudy (anonymous):

Operations of functions, its quite simple but I don't understand it. Find (a): (f+g)(3) 1) f(x)=x+3 g(x)=x^2 How do I do this?

OpenStudy (anonymous):

@mathstudent55

OpenStudy (aum):

f(x)=x+3 g(x)=x^2 (f + g)(x) = f(x) + g(x) = x^2 + x + 3 (f + g)(3) = ?